OpenStudy (anonymous):

find all solutions to the equation sin^2(2x)=1/4

OpenStudy (anonymous):

you're going to be writing for a while.. (there's infinite solutions)

OpenStudy (anonymous):

[0,2pi]

OpenStudy (anonymous):

x = 1/12 (6 pi n+-pi)

OpenStudy (anonymous):

\[\Large \sin^{2}(2x)=\frac{1}{4}\] \[\Large \sin^{2}(2x)=\sin^{2}\frac{\pi}{6}\] Now we know that for- \[\Large \sin^{2}\theta=\sin^{2}\alpha\] \[\Large \theta=n\pi \pm\alpha \] substitute theta and alpha.